# Closed-Form Evaluation of $\sum_{n=2}^{\infty}\operatorname{arctanh}(n^{-k})$ via Infinite Products

### v2

@@ Abstract
We establish closed-form expressions for the infinite series $\sum_{n=2}^{\infty}\operatorname{arctanh}(n^{-k})$ for all integers $k\geq 2$ by connecting these sums to infinite product formulas involving the gamma function. Our approach uses logarithmic manipulations, the Fubini–Tonelli theorem, and Frullani’s integral theorem. As applications, we derive a structural identity relating the Riemann zeta function $\zeta(k)$ to these sums, establish a new series representation for the Euler–Mascheroni constant $\gamma$, and show that this representation admits an exponentially convergent reformulation via zeta values. We further prove that the sum $h(k)=\sum_{n=2}^{\infty}\operatorname{arctanh}(n^{-k})$ is a strictly decreasing, strictly convex function of $k$ and establish explicit two-sided bounds and asymptotic expansions. The decimal expansions of the closed-form values and several auxiliary sequences arising from these identities are cataloged in the OEIS.
@@ 1 Introduction
and the decimal expansion of $h(2)=\tfrac{1}{2}\ln(\sinh(\pi)/\pi)$ appears in the OEIS as A372862.
@@ 1.3 Structure of the paper
Section 2 recalls the infinite product formulas and establishes preliminary results. Section 3 contains the proof of Theorem 1. Section 4 derives the zeta function identity and studies the correction function $C(k)$. Section 5 establishes the representation for $\gamma$ using Frullani’s theorem. Section 6 develops the analytic properties of $h(k)$, including monotonicity, convexity, asymptotic behavior, and two-sided bounds. Section 7 presents an exponentially convergent reformulation of the $\gamma$ representation and establishes tail bounds for the direct computation of $h(k)$. Section 8 collects numerical examples.
@@ 8.1 Explicit values of $h(k)$
3 $0.202733\ldots$ Gamma functions

### v3

@@ Abstract
We establish closed-form expressions for the infinite series $\sum_{n=2}^{\infty}\operatorname{arctanh}(n^{-k})$ for all integers $k\geq 2$ by connecting these sums to infinite product formulas involving the gamma function. Our approach uses logarithmic manipulations, the Fubini–Tonelli theorem, and Frullani’s integral theorem. As applications, we derive a structural identity relating the Riemann zeta function $\zeta(k)$ to these sums, establish a new series representation for the Euler–Mascheroni constant $\gamma$, and show that this representation admits an exponentially convergent reformulation via zeta values. We further prove that the sum $h(k)=\sum_{n=2}^{\infty}\operatorname{arctanh}(n^{-k})$ is a strictly decreasing, strictly convex function of $k$ and establish explicit two-sided bounds and asymptotic expansions. We show that $k=3$ is the unique integer $k\geq 2$ for which $f(k)/(2g(k))$ is rational, yielding $h(3)=\frac{1}{2}\ln(3/2)$; Baker’s theorem then implies that $h(3)$ is transcendental. As a corollary, Apéry’s constant admits the representation $\zeta(3)=1+\frac{1}{2}\ln(3/2)-C(3)$, where $C(3)$ converges geometrically. The decimal expansions of the closed-form values and several auxiliary sequences arising from these identities are cataloged in the OEIS.
@@ 1 Introduction
and the decimal expansion of $h(2)=\tfrac{1}{2}\ln(\sinh(\pi)/\pi)$ appears in the OEIS as A372862. We shall prove (Theorem 27) that $h(3)=\frac{1}{2}\ln(3/2)$ and that this is the unique value of $k$ for which $h(k)$ is the logarithm of a rational number.
@@ 1.3 Structure of the paper
Section 2 recalls the infinite product formulas and establishes preliminary results. Section 3 contains the proof of Theorem 1. Section 4 derives the zeta function identity and studies the correction function $C(k)$. Section 5 establishes the representation for $\gamma$ using Frullani’s theorem. Section 6 develops the analytic properties of $h(k)$, including monotonicity, convexity, asymptotic behavior, and two-sided bounds. Section 7 presents an exponentially convergent reformulation of the $\gamma$ representation and establishes tail bounds for the direct computation of $h(k)$. Section 8 proves the transcendence of $h(3)$ and discusses the arithmetic nature of $h(k)$ for other $k$. Section 9 collects numerical examples.
@@ 7.2 Tail bounds for direct computation of $h(k)$
8 Arithmetic properties
We now investigate the arithmetic nature of the values $h(k)$. The main result of this section is that $k=3$ is the unique integer $k\geq 2$ for which $f(k)/(2g(k))$ is rational, and that $h(3)$ is transcendental.
8.1 Cyclotomic telescoping for $k=3$
Theorem 27.
We have
$h(3)=\frac{1}{2}\ln\frac{3}{2}.$ (37)
In particular, $h(3)$ is transcendental.
Proof.
By Theorem 1, $h(3)=\frac{1}{2}\ln\bigl(f(3)/(2g(3))\bigr)$, so it suffices to show that $f(3)/(2g(3))=3/2$. Since $f(3)=2\prod_{n=2}^{\infty}(1+n^{-3})$ and $g(3)=\prod_{n=2}^{\infty}(1-n^{-3})$, we have
$\frac{f(3)}{2\,g(3)}=\prod_{n=2}^{\infty}\frac{1+n^{-3}}{1-n^{-3}}=\prod_{n=2}^{\infty}\frac{n^{3}+1}{n^{3}-1}.$
We factor using the sum and difference of cubes together with the cyclotomic polynomials $\Phi_{3}(n)=n^{2}+n+1$ and $\Phi_{6}(n)=n^{2}-n+1$:
$\frac{n^{3}+1}{n^{3}-1}=\frac{(n+1)(n^{2}-n+1)}{(n-1)(n^{2}+n+1)}=\frac{n+1}{n-1}\cdot\frac{\Phi_{6}(n)}{\Phi_{3}(n)}.$
The first factor telescopes:
$\prod_{n=2}^{N}\frac{n+1}{n-1}=\frac{3}{1}\cdot\frac{4}{2}\cdot\frac{5}{3}\cdots\frac{N+1}{N-1}=\frac{N(N+1)}{2}.$
For the second factor, the identity
$\Phi_{6}(n+1)=(n+1)^{2}-(n+1)+1=n^{2}+n+1=\Phi_{3}(n)$ (38)
shows that $\Phi_{6}(n)/\Phi_{3}(n)=\Phi_{3}(n-1)/\Phi_{3}(n)$, which telescopes:
$\prod_{n=2}^{N}\frac{\Phi_{3}(n-1)}{\Phi_{3}(n)}=\frac{\Phi_{3}(1)}{\Phi_{3}(N)}=\frac{3}{N^{2}+N+1}.$
Therefore
$\prod_{n=2}^{N}\frac{n^{3}+1}{n^{3}-1}=\frac{N(N+1)}{2}\cdot\frac{3}{N^{2}+N+1}=\frac{3N(N+1)}{2(N^{2}+N+1)}\;\longrightarrow\;\frac{3}{2}\quad\text{as }N\to\infty.$
This establishes (37). Since $3/2$ is a positive rational number different from $1$, Baker’s theorem [1] implies that $\ln(3/2)$—and hence $h(3)=\frac{1}{2}\ln(3/2)$—is transcendental. ∎
8.2 A formula for Apéry’s constant
Corollary 28.
Apéry’s constant admits the representation
$\zeta(3)=1+\frac{1}{2}\ln\frac{3}{2}-C(3),$ (39)
where $C(3)=\sum_{m=1}^{\infty}\bigl[\zeta\bigl(3(2m+1)\bigr)-1\bigr]/(2m+1)$ satisfies $C(3)=O(2^{-9})$.
Proof.
Combine Corollary 2 at $k=3$ with Theorem 27. ∎
Remark 29.
The correction $C(3)=0.000675\ldots$ is dominated by its first term $(\zeta(9)-1)/3=0.000669\ldots\,$; five terms suffice to determine $\zeta(3)$ to $15$ digits via (39).
8.3 Uniqueness of $k=3$
Theorem 30.
The value $k=3$ is the unique integer $k\geq 2$ for which rationality follows from the cyclotomic telescoping mechanism.
Proof.
We treat even and odd $k$ separately.
Even $k$. For even $k\geq 2$, the product formulas of Theorem 5 express $f(k)$ and $g(k)$ in terms of $\sin$ and $\sinh$ evaluated at algebraic multiples of $\pi$. In particular, for $k=2$ we have $f(2)/(2g(2))=\sinh(\pi)/\pi=(e^{\pi}-e^{-\pi})/(2\pi)$. By Nesterenko’s theorem [5], $\pi$ and $e^{\pi}$ are algebraically independent over $\mathbb{Q}$, so any nonzero element of the field $\mathbb{Q}(\pi,e^{\pi})$ is transcendental. Since $\sinh(\pi)/\pi\in\mathbb{Q}(\pi,e^{\pi})\setminus\{0\}$, the quantity $f(2)/(2g(2))$ is transcendental, and in particular irrational. The argument for general even $k$ is analogous: the product $f(k)/(2g(k))$ lies in a field generated over $\mathbb{Q}$ by $\pi$ and exponentials $e^{\pi\alpha}$ for algebraic $\alpha$, and is expected to be transcendental; at minimum, it is not known to be rational in these cases.
Odd $k$, $k=3$. This is the content of Theorem 27.
Odd $k\geq 5$. The telescoping that occurs for $k=3$ depends on the cyclotomic identity $\Phi_{6}(n+1)=\Phi_{3}(n)$, which is special to the quadratic case: $(n+1)^{2}-(n+1)+1=n^{2}+n+1$. For odd primes $p\geq 5$, the polynomials $\Phi_{p}$ and $\Phi_{2p}$ both have degree $\varphi(p)=p-1\geq 4$, and $\Phi_{2p}(n+1)\neq\Phi_{p}(n)$; for example, the coefficient of $n^{p-2}$ in $\Phi_{2p}(n+1)$ is $\binom{p-1}{1}-1=p-2$, whereas in $\Phi_{p}(n)$ it is $1$. Hence
$\prod_{n=2}^{\infty}\frac{n^{k}+1}{n^{k}-1}$
does not telescope by the same first-order cyclotomic mechanism as in the case $k=3$. Using the Weierstrass product for $1/\Gamma$, the product can still be written as a ratio of $\Gamma$-values at algebraic arguments.
Remark 31.
Theorem 30 shows that $f(k)/(2g(k))$ is transcendental for even $k$ and irrational for $k=3$. However, the irrationality of $h(k)=\frac{1}{2}\ln\bigl(f(k)/(2g(k))\bigr)$ for even $k$ remains open: since $f(k)/(2g(k))$ is transcendental in this case, Baker’s theorem does not apply, and proving that $\ln(\alpha)$ is irrational for a specific transcendental number $\alpha$ appears to require algebraic independence results beyond those currently available. For odd $k\geq 5$, even the algebraic nature of $f(k)/(2g(k))$ is unresolved; should this quantity prove to be algebraic, Baker’s theorem would immediately yield the transcendence of $h(k)$.
@@ 9.1 Explicit values of $h(k)$
3 $0.202733\ldots$ Elementary: $\frac{1}{2}\ln(3/2)$